Opened 11 years ago

Last modified 2 years ago

#371 closed flight processing

EUFAR10/01, flight day 238b/2010, Lake Balaton Peninsula — at Version 20

Reported by: iopa Owned by: knpa
Priority: alpha 4 high Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anch)

Data location: ~arsf/workspace/EU10_03-2010_238b_Lake_Balaton_Peninsula

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Mapping of species composition and stand health of shore vegetation, mapping of shore elevation and inundation structure.

This flight will also be used for EUFAR10-07.
Scientific Objective: High accuracy land use/land cover mapping, forest type mapping and invasive shrub species location

Priority: a4h

PI: A. Zlinszky

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

Sensors:

  • Eagle
  • Hawk
  • Leica LiDAR (04/03/2011)
  • Digital Photography (04/03/2011)

Change History (20)

comment:1 Changed 11 years ago by iopa

  • Description modified (diff)

comment:2 Changed 11 years ago by iopa

  • Description modified (diff)
  • Priority changed from alpha 5 to alpha 4 high

comment:3 Changed 11 years ago by iopa

The applanix and IPAS for this project have been copied from 238a. The processing need only be done once.

comment:4 Changed 11 years ago by iopa

Note; Eagle file VNIR238b-10-3.raw is still incomplete and needs to be recovered when ARSF Ops notify us that it is on thelma.

comment:5 Changed 11 years ago by knpa

  • Description modified (diff)

Edited scientific objective.

comment:6 Changed 11 years ago by knpa

  • Description modified (diff)

This flight will also be used for EUFAR10-07.

comment:7 Changed 11 years ago by iopa

From an email by Phil Goy on 19/10/2010,

Eagle file VNIR238b-10-3.raw is unrecoverable and we will have to process only the partial
one we already have.

The header for that still needs to be fixed to match the partial raw.

comment:8 Changed 11 years ago by iopa

The header has been fixed and the line is processable as any other, BUT

(for eagle line 3)
azspec_args_extra needs to contain (at least):
azspec_args-extra = -sdr eagle/darkFrames.bil -sdh eagle/darkFrames.bil.hdr

Those dark frames have been extracted from VNIR238b-10-1.raw, it has been tested as far as azspec and seems to work. If any problems arise in next stages for this line, keep this in mind.

comment:9 Changed 11 years ago by iopa

It turned out, as Phil and James told us when they visited, the above file was successfully recovered and then uploaded on thelma. After downloading and checking it, the file is complete and processable.

None of the above 'special care' will be necessary.

comment:10 Changed 11 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:11 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Taking LIDAR.

comment:12 Changed 10 years ago by emca

Processed the camera imagery from raw to tiff

comment:13 Changed 10 years ago by knpa

Processed LIDAR. Images match-up well horizontally and vertically.

comment:14 Changed 10 years ago by knpa

Delivery created and ready for checking at: ~airborne/workspace/EUFAR10_01-2010_238b_Lake_Balaton_Peninsula/delivery/20110221/EUFAR10_01

comment:15 Changed 10 years ago by iopa

Checking LiDAR.

comment:16 Changed 10 years ago by iopa

LiDAR delivery check;

no errors found, data ready for delivery.

comment:17 Changed 10 years ago by emca

Processed camera data - ready for delivery check at
~airborne/workspace/EUFAR10_01-2010_238b_Lake_Balaton_Peninsula/delivery/20110302/EUFAR10_01

comment:18 Changed 10 years ago by anch

Starting camera delivery check

comment:19 Changed 10 years ago by anch

Camera delivery check complete.

All good and ready to deliver.

comment:20 Changed 10 years ago by anch

  • Description modified (diff)

Camera and LiDAR sent to PI

Note: See TracTickets for help on using tickets.